# Ontology-driven Reinforcement Learning for Personalized Student Support

### v1

@@ I Introduction
Fitting with AI’s ability to augment human intelligence, applying AI to the task of personalized educational support is a natural conclusion to address resource limitations. AI is easily integrated into virtual educational systems such as serious games and intelligent tutoring systems [6, 7]. With AI and virtual educational systems, several other problems can be addressed at the same time. First, students can be given active, timely feedback, regardless of whether they are learning in the classroom, at an after-school study group, or at home. Second, the AI can actively prompt and engage with students, helping them to overcome situations where they may fail to even formulate a proper question [8]. Finally, AI can easily be integrated with a lesson plan or educational database to answer questions or address any misunderstandings that a student may have without the need for instructor intervention. With AI as a study aid, well-performing students can cover their areas of difficulty while poor-performing students can catch up without excessive instructor intervention.
@@
To integrate an AI into a virtual educational system, it is first necessary to create a knowledge database or lesson plan that the AI can pull from. Such a database is necessary to both track student performance on different topics and to allow for human intelligence to organize knowledge for the AI. With that framework, the job of the AI is purely content delivery, and factual errors are uncommon. For our proposed method, this knowledge database is created with an ontology; a type of data structure used to model and organize knowledge [9]. With their structure, ontologies are a powerful framework, usable for organizing educational content in a hierarchical manner while creating a format that can be easily connected with an AI. In literature, ontologies have shown effectiveness for student modelling for recommender systems [10] and intelligent tutoring systems [11], simplifying lesson planning while accounting for educational standards [12], and for representing general educational content [13].
With an ontology representing the domain knowledge that a virtual educational system intends to teach, the remaining components of the proposed system serve as the ”brain” of the AI. In this case, we adapt multi-agent reinforcement learning (RL) [14] to act as an adaptive controller that can adjust its behavior based on observations of students. This dynamic adaptability both guarantees that the AI will eventually perform optimally [15], and allows the AI system to account for shifting trends in student behavior. In our prior work, we developed and tested a multi-agent RL approach using experience sharing to improve agent training performance [16]. We verified the positive performance impact of this method [17], and have successfully applied this method within an educational serious game, showing a positive impact on student performance [17].
However, our prior work applied the multi-agent reinforcement learning to a single specific case, dealing with a small subset of topics within a single domain. It wasn’t straightforward to add new topics to the system, and furthermore, the system wasn’t generalized for wider applications. By adapting this method to the new ontology-based structure, multiple RL agents can be connected to various topics within the ontology to each provide specialized support on their given topic. With this structure, the ontology can easily be extended like a modular system, adding or removing RL agents as needed. By adapting our prior experience sharing method, newly-added agents can boost their performance using shared experience, giving them a jump-start at the beginning of training to avoid a period of poor performance at the start. And ultimately, the result is a modular, easily-extensible student support system that can function in various virtual educational systems on various topics.
The rest of this paper is organized as follows: Section II details our proposed method on a theoretical level, discussing the components of our proposed system with ontologies in Section II-A, data collection and formatting in Section II-B, reinforcement learning methodology in Section II-D, and personalized support generation in Section II-E. Finally, Section III provides conclusions and future directions for this research.
@@ II-A Ontologies for Education
For this application, ontologies were chosen due to their past successes in educational applications, including student modelling [10, 11], and educational content organization [13, 12]. Furthermore, the semantic connections used allow for a more meaningful structure, heavily informing the lesson plan and educational content presentation of any connected system. Finally, the organization also supports the proposed system, allowing for sensible organization of both the reinforcement learning agents used as the ”brain” of the system, and of the topic-specific support databases required to generate personalized student assistance.
The organization of system’s ontology is very important to the overall configuration of the system. The ontology determines the order of operations that the system follows when teaching a student, as well as the topics that are assigned reinforcement learning agents. Since agents each specialize in a specific topic, agent-augmented nodes in the ontology must be granular enough where a specialized agent can provide relevant and useful assistance, while still being broad enough to not require an exponential number of RL agents. In the case of Figure 2, we might assign an RL agent to the nodes for ”properties of a function”, ”linear function”, ”quadratic function”, and ”exponential function” to provide students with support on those topics. Then, any topic under the selected topics can be used to inform tutoring and generate personalized recommendations, as will be discussed in a later section.
@@
From a purely educational approach, the organization of these topics is similar to how an educator would configure a lesson plan. Thus, the integrated system can use the educational ontology as a point of reference for how the educational content in the game should be organized. This data structure also encompasses the necessary information to teach the enclosed content, and, when applicable, the reinforcement learning agents assigned to select topics.
@@ II-B Data Collection and Storage
With the ontology defined, the second step in system creation is student data collection. Given an ontology $O^{*}$ and its set of concepts $C^{*}$, the system collects student data for each ontology concept $c_{i}\in C^{*}$. By matching student data to each topic in the ontology, the system has a solid understanding of a student’s performance in each of the topics. We formalize this in the proposed system by stating that each concept in the ontology stores a student data vector, $\mathbf{x_{c_{i}}}$, such that this vector contains a set of generic performance metrics that can indicate a student’s performance in any given topic.
The specific performance metrics used depend on implementation, but should be general-purpose such that any topic can use the same metrics. This shared form for every node is important to the reinforcement learning, as will be discussed in Section II-D. To give some examples, $\mathbf{x_{c_{i}}}$ could be defined as $\mathbf{x_{c_{i}}}=\begin{pmatrix}a&b&c\end{pmatrix}$, where $a$ is a measure of competency, such as a quiz score, $b$ is a measure of timing, such as a time-to-completion measurement, and $c$ is a numerical indicator of engagement, such as number of interactions made or time spent idle. These metrics are purposefully left in a generic format to allow the proposed system to fit a wide range of topics or situations.
@@ II-C Data Transformation
As mentioned, each data vector $\mathbf{x_{c_{i}}}$ has a shared form for every topic, and is meant to represent a general-purpose vector (referred to here as $\mathbf{x}$) of student performance metrics that can be used for any educational topic or scenario. However, the inherent diversity in data collection methods also requires the proposed system to account for educational platforms that may collect data in varied forms. Therefor, it is necessary for the proposed system to transform incoming data from some unknown set of collected metrics to a known list of accepted metrics. Additionally, to improve reinforcement learning agent performance, it is also necessary to normalize incoming data within a specified range as part of the data transformation process.
Connecting with other sections of the proposed system architecture, data transformation begins with a semantic representation of the target data metrics in $\mathbf{x}$. With this representation, it is necessary to assign a set of descriptors to each metric in $\mathbf{x}$. Then, if the system has a new vector $\mathbf{y_{c_{i}}}$ of input data with format $\mathbf{y}$, containing some varied set of metrics each also with descriptors attached, $\mathbf{y}$ can be connected within the data transformation ontology to $\mathbf{x}$. With connections formed, let $f:\mathbf{y}\xrightarrow{}\mathbf{x}$ define a set of transformation functions that map each element in $\mathbf{y}$ to an element in $\mathbf{x}$, performing necessary data transformation and normalization to translate $\mathbf{y}$ into $\mathbf{x}$. Figure 3 shows an example of how this data transformation can be formalized. For example, if $y_{1}$ ranges in $[0,10]$ and $y_{3}$ ranges in $[0,20]$, a possible simple function for $f_{1}(y_{1},y_{3})$ is shown in Equation 1. In this equation, both inputs are simply divided by their maximums, added together, and then divided again by 2 to get an average of the two. This both compresses the two data points down to a single expected data point and normalizes $x_{1}$ to between $[0,1]$. However, this function is only a simplified example, and functions in the proposed system could take into account more or less data points, consider more complex statistical methods for data transformation, or even be completed by a non-linear approximator such as a neural network.
$f_{1}(y_{1},y_{3})=(\frac{y_{1}}{10}+\frac{y_{3}}{40})/2=x_{1}$ (1)
@@ II-D Multi-agent Reinforcement Learning
Reinforcement learning provides a natural fit for the problem of personalized student assistance. Given the established data vector, $\mathbf{x_{c_{i}}}$, that stores student data relative to a given topic $c_{i}\in C^{*}$ in our example ontology $O^{*}$, this data can represent a good environmental state for an RL agent. The iterative online learning that reinforcement learning agents provide means that system developers do not have to encode complex decision logic to provide personalized support, instead relying on the RL agents to naturally learn optimal behavior. Furthermore, since reinforcement learning is an ongoing, online process, shifting trends in student preferences or slight changes in lesson planning can be adapted to, ensuring that students always receive appropriate personalized assistance.
@@
    To exemplify this configuration, a possible action vector $\mathbf{a}$ might be defined as $\mathbf{a}=\begin{pmatrix}a_{1}&a_{2}&a_{3}\end{pmatrix}$, where $a_{1}$ is a weight for visual content, such as images; $a_{2}$ is a weight for example problems, and; $a_{3}$ is a weight for video tutorials that require the student to follow along. Thus, if $\mathbf{a}=\begin{pmatrix}1.0&0.0&0.0\end{pmatrix}$, the system is recommending personalized assistance that heavily prioritizes images and visual examples, while deprioritizing example problems and video tutorials.
@@ II-E Generating Personalized Assistance
III Conclusions

### v2

@@ I Introduction
Fitting with AI’s ability to augment human intelligence, AI is easily integrated into virtual educational systems such as serious games and intelligent tutoring systems [6, 7]. With AI and virtual educational systems, students can be given active, timely feedback, regardless of whether they are learning in the classroom, at an after-school study group, or at home. Furthermore, the AI can actively prompt and engage with students, helping them to overcome situations where they may fail to even formulate a proper question [8]. Finally, AI can easily be integrated with a lesson plan or educational database to answer questions or address any misunderstandings that a student may have without the need for instructor intervention. With AI as a study aid, well-performing students can cover their areas of difficulty while poor-performing students can catch up without excessive instructor intervention.
AI integration in virtual educational systems first require some knowledge database or lesson plan to guide education, track students’ progress, and allow for human input to ensure accurate lessons. For our proposed method, this knowledge database is created with an ontology; a type of data structure used to model and organize knowledge [9]. With their structure, ontologies are a powerful framework, usable for organizing educational content in a hierarchical manner while creating a format that can be easily connected with an AI. In literature, ontologies have shown effectiveness for student modelling for recommender systems [10] and intelligent tut-
oring systems [11], simplifying lesson planning while accounting for educational standards [12], and for representing general educational content [13].
@@
With an ontology representing the domain knowledge that a virtual educational system intends to teach, the remaining components of the proposed system serve as the ”brain” of the AI. We adapt multi-agent reinforcement learning (RL) [14] to act as an adaptive AI tutor. This dynamic adaptability both guarantees that the AI will eventually perform optimally [15], and allows the AI system to account for shifting trends in student behavior. In our prior work, we developed and tested a multi-agent RL approach using experience sharing to improve agent training performance [16]. We verified the positive performance impact of this method [17], and have successfully applied this method within an educational serious game, showing a positive impact on student performance [17].
However, our prior work applied the multi-agent reinforcement learning to a single case, dealing with a small subset of topics within a single domain. It wasn’t straightforward to add new topics to the system, and furthermore, the system wasn’t generalized for wider applications. With the proposed work, we aim to extend the system to support multiple connected RL agents using the ontology structure for organization. By adapting our prior experience sharing method, newly-added agents can boost their performance using shared experience, giving them a jump-start at the beginning of training to avoid a period of poor performance at the start. The result is a modular, easily-extensible student support system that can function in various virtual educational systems on various topics.
@@ II-A Ontologies for Education
For this application, ontologies were chosen due to their past successes in educational applications, including student modelling [10, 11], and educational content organization [13, 12]. Furthermore, the semantic connections used allow for a more meaningful structure, informing the lesson plan and educational content presentation of any connected system. In other words, topics can be organized similarly to a lesson plan, helping to inform both AI assistance and the overall order of content presentation.
The ontology determines the order of operations that the system follows when teaching a student, as well as the topics that are assigned reinforcement learning agents. As the RL agents are meant to specialize in topics, agent-augmented nodes in the ontology must be granular enough where a specialized agent can provide relevant and useful assistance, while still being broad enough to not require a massive number of RL agents. In the case of Figure 2, we might assign an RL agent to the sub-classes ”linear function”, ”quadratic function”, and ”exponential function”, as well as one for the properties of domain and range. Then, any topic under the selected topics can be used to inform tutoring and generate personalized recommendations, as will be discussed in a later section.
@@ II-B Data Collection and Storage
With the ontology defined, the second step in system creation is student data collection. Given an ontology $O^{*}$ and its set of concepts $C^{*}$, the system collects student data for each concept $c_{i}\in C^{*}$. By matching student data to each topic in the ontology, the system has a solid understanding of a student’s performance in each of the topics. We formalize this in the proposed system by stating that each concept in the ontology stores a student data vector, $\mathbf{x_{c_{i}}}$, such that this vector contains a set of generic performance metrics that can indicate a student’s performance in any given topic.
The specific performance metrics used depend on implementation, but should be general-purpose such that any topic can use the same metrics. This shared form for every node is important to the reinforcement learning, as will be discussed in Section II-D. To give some examples, $\mathbf{x_{c_{i}}}$ could be defined as $\mathbf{x_{c_{i}}}=\begin{pmatrix}a&b&c\end{pmatrix}$, where $a$ is a measure of competency, such as a quiz score, $b$ is a measure of timing, such as a time-to-completion measurement, and $c$ is a numerical indicator of engagement, such as number of interactions made or time spent idle. Ultimately, the data points chosen depend on implementation.
@@ II-C Data Transformation
As mentioned, each data vector $\mathbf{x_{c_{i}}}$ has a shared form for every topic, and is meant to represent a general-purpose vector (referred to here as $\mathbf{x}$) of student performance metrics that can be used for any educational topic or scenario. However, the inherent diversity in data collection methods also requires the proposed system to account for variation in collected student data. Therefore, it is necessary for the proposed system to transform incoming data from some unknown set of collected metrics to a known list of accepted metrics.
Connecting with other sections of the proposed system architecture, data transformation begins with a semantic representation of the target data metrics in $\mathbf{x}$. With this representation, it is necessary to assign a set of descriptors to each metric in $\mathbf{x}$. Then, if the system has a new vector $\mathbf{y_{c_{i}}}$ of input data with format $\mathbf{y}$, containing some varied set of metrics each also with descriptors attached, $\mathbf{y}$ can be connected within the data transformation ontology to $\mathbf{x}$. With connections formed, let $f:\mathbf{y}\xrightarrow{}\mathbf{x}$ define a set of transformation functions that map each element in $\mathbf{y}$ to an element in $\mathbf{x}$, performing necessary data transformation and normalization to translate $\mathbf{y}$ into $\mathbf{x}$. Figure 3 shows an example of how this data transformation can be formalized. Section III shows a more in-depth example of this data transformation principle.
@@ II-D Multi-agent Reinforcement Learning
Reinforcement learning provides a natural fit for the problem of personalized student assistance. Given the established data vector, $\mathbf{x_{c_{i}}}$, that stores student data relative to a given topic $c_{i}\in C^{*}$ in our example ontology $O^{*}$, this data can represent an environmental state for an RL agent. The iterative online learning that reinforcement learning agents provide means that system developers do not have to encode complex decision logic to provide personalized support, instead relying on the RL agents to naturally learn optimal behavior. Furthermore, since reinforcement learning is an ongoing, online process, shifting trends in student preferences or slight changes in lesson planning can be adapted to, ensuring that students always receive appropriate personalized assistance.
@@
    Section III gives an in-depth example of how an action vector might be defined using properties of personalized student assistance.
@@ II-E Generating Personalized Assistance
III Example Implementation
Given the system definition, we now provide a brief example of how such a system might be implemented within some educational software. In this example educational software, we can assume that the system provides educational material to students and intermittently administers quizzes to gauge each student’s content knowledge. With that in mind, we first must consider the data collected by such a system, defining the data input vector, $\mathbf{y}=\{y_{1},y_{2},y_{3},y_{4},y_{5}\}$ with the following values:
  • $y_{1}$: A given student’s percentage of correct answers on a quiz in range $[0,1]$
  • $y_{2}$: The total amount of time spent on a given section, in seconds
  • $y_{3}$: The amount of time spent engaging with educational content (ex. reading material, answering questions), in seconds
  • $y_{4}$: The total number of button presses or other interactions the student has made within a given section
  • $y_{5}$: An emotional state estimated from webcam images of the student as they play, with 00 indicating negative emotions, $1$ indicating neutral, and $2$ indicating positive
With those values defined, we can now determine the mapping functions to translate the given data into the system’s input vector, $\mathbf{x}=\{x_{1},x_{2},x_{3}\}$, where:
  • $x_{1}$: A numerical score in range $[0,1]$ indicating a student’s competency on the provided content
  • $x_{2}$: A numerical score in range $[0,1]$ indicating a student’s engagement with the provided content
  • $x_{3}$: A numerical score indicating a student’s current emotional state in range $[0,1]$, where 00 indicates that frustration might be present, while $1$ indicates a positive emotional state with no measured issues
And then the mapping functions can determine the relationship between the actual gathered data and the data vector sent in to the RL system. Quiz scores ($y_{1}$) can easily map directly into the value for competence ($x_{1}$), as shown in Equation 1. Engagement ($x_{2}$) can be determined based on the student’s number of interactions with the system ($y_{4}$) and how they compare to some maximum value ($y_{4}^{max}$) determined through testing. As shown in Equation 2, engagement can also take into account the percentage of time a student spends on educational content ($y_{3}$) compared to their total time in the game ($y_{2}$) to provide further insight into their engagement with the provided material. Finally, as shown in Equation 3, emotional state can use a scaled version of the emotional measurements ($y_{5}$), but can be adjusted if the user is spending an above-average amount of time (with $y_{2}^{avg}$ indicating average time) on a given section.
$x_{1}=y_{1}$ (1)
$x_{2}=(\frac{y_{4}}{y_{4}^{max}}+\frac{y_{3}}{y_{2}})/2$ (2)
$x_{3}=\begin{cases}\frac{y_{5}}{2}&y_{2}\leq y_{2}^{avg}\\ \max{(\frac{y_{5}}{2}-\frac{y_{2}-y_{2}^{avg}}{y_{2}^{avg}},0)}&y_{2}>y_{2}^{avg}\end{cases}$ (3)
Given these mapping functions, we aim to demonstrate that the proposed system can be flexible, with a variety of acceptable data collection methods and data inputs all mapping to a standard format so that different systems can use similar RL architectures. An actual implementation could extend this basic system to include significantly more complex mapping functions, such as neural networks, machine learning methods, or statistical analyses. With $\mathbf{x}$ fully mapped, the system can use it as the RL state to generate the action vector. In our example system, we define the action vector $\mathbf{a^{*}}=\{a_{1},a_{2},a_{3},a_{4},a_{5}\}$, where:
  • $a_{1}$ is a weight for how heavily the generated assistance should rely on visuals.
  • $a_{2}$ is a weight for the inclusion examples, such as videos or step-by-step guides, that should be included in the generated assistance.
  • $a_{3}$ is a weight for the inclusion of practice problems.
  • $a_{4}$ is a weight for how much guidance the provided assistance should give to players. For example, a low value here would mean that the player simply receives educational material and must learn, while a high value here would mean that the system gives significantly more detailed guidance on what the player should study.
  • $a_{5}$ is a weight for how encouraging and emotionally-driven the dialogue given to the player should be.
All values are essentially treated as weights in the range $[0,1]$ and interpreted by the personalized assistance generation module which might combine rule-based systems with large language models. For example, with the weights given above, a rule-based system could first interpret $a_{1},a_{2},$ and $a_{3}$, which are the weights for visuals, examples, and practice problems. We could then take an expert-created set of study materials and assign threshold values to each piece of content. Some images might only be shown to students who were assigned an $a_{1}$ value of $0.9$ or higher, with similar rules in place for example problems and practice problems. Meanwhile, $a_{4}$ and $a_{5}$ are more abstract, so a large language model could be used to dynamically regenerate existing educational content to add in more or less guidance based on $a_{4}$, and more or less encouragement based on $a_{5}$. With this simple example, students would see highly varied study material based on the RL system’s decision-making, and student feedback on the usefulness of that material could then feed back into the RL to inform decisions for future students.
IV Conclusions